# LIMOncello: Revisited IKFoM on the SGal(3) Manifold for Fast LiDAR–Inertial Odometry

### v1

@@ I Introduction
Light detection and ranging (LiDAR) sensors complements vision in such setups by providing long-range, illumination-invariant, and metrically accurate depth, which translates into reliable odometry and mapping across diverse environments. However, LiDAR introduces distinct challenges: i) processing a massive, continuous stream of data in a timely and efficient manner, ii) overcome the scarcity of reliable geometric features in structurally poor environments, and iii) correcting the scan deformation caused by sensor movement during data capture [21, 20].
@@
To address these issues, LiDAR–Inertial Odometry (LIO) combines LiDAR with high-rate inertial measurements. High-rate Inertial Measurement Unit (IMU) measurements enable per-point deskewing and cover the latency between LiDAR frames, while LiDAR provides metric structure to correct the robot’s pose. Most LIO systems are organized into a front-end and a back-end [1]. The front-end maintains a history map and performs fast data association. This process is based on classical point-cloud registration, which works by finding the nearest neighbors for each incoming point. These neighbors are then used to fit a simple local model (typically a plane or a line) and form the residuals for optimization.
@@
The back-end is responsible for state optimization, with the two common families being graph-based smoothing and filtering. Under limited onboard processing, filtering based on the Extended Kalman Filter and its iterated error-state (IESKF) variant is often preferred [11], as exemplified by FAST-LIO1 [21] and FAST-LIO2 works [20].
@@
A key step forward was the IKFoM [11] framework, which formalized the IESKF on differentiable manifolds and packaged it into a practical toolkit. This gave a canonical discrete-time representation and clean separation between manifold operations and system-specific models.
@@
With a robust and reusable back-end in place, the community has focused effort on the front-end problems of features and computation: designing faster spatial data structures and residual evaluations that sustain real-time rates without sacrificing accuracy. The community has improved classical k-d tree map structures with incremental variants such as the ikd-Tree [3] and sparse voxel representations like iVox [1], enabling efficient real-time updates and queries. Building on these advances, recent methods have shifted from point-to-map registration to point-to-feature formulations, where points are aligned with local geometric primitives (e.g., planes or edges) rather than individual map points [19, 22].
@@
Alongside discrete-time pipelines, continuous-time motion models have gained attraction. By representing the trajectory with piecewise functions and evaluating residuals at each point timestamp, these methods improve deskewing and alignment during fast motion or asynchronous sensing. Recent systems such as DLIO [5] adopt point-wise continuous-time motion compensation, and others like RESPLE [4] show that investing in state representation and motion modeling can yield competitive accuracy while preserving real-time performance.
@@
Most LIO frameworks rely on what is often referred to as a compound [11] or bundle manifold [17] representation of the state. In these formulations, the system’s pose, and velocity are defined on separate sub-manifolds—usually $\mathrm{SO}(3)\times\mathbb{R}^{6}$. While this decomposition simplifies the implementation, it also neglects certain kinematic couplings between translational and rotational motion, leading to an approximation that may lose accuracy under highly dynamic motion or under sustained low-observability conditions, e.g., near-zero acceleration that reduces IMU observability, or degenerate LiDAR environments such as long corridors.
@@
Beyond Euclidean and compound representations, several Lie-group manifolds have been adopted in robotics. $\mathrm{SE}(3)$ is widely used for pose estimation, while $\mathrm{SE}_{2}(3)$ became standard in InEKF-based methods [2], with extensions such as $\mathrm{SE}_{N}(3)$ [23] and $\mathrm{SGal}(3)$ explored more recently in equivariant filtering [6]. However, their use in tightly coupled LIO-SLAM systems remains limited.
@@
  • Efficient spatial data structure: We replace the widely used ikd-Tree with a refactored, lightweight i-Octree implementation [24], achieving faster incremental updates and lower computational overhead by avoiding costly tree rebalancing.
  • Modular and extensible implementation: LIMOncello is implemented using the manif library [7], resulting in a clean, modular codebase that is easily extensible and publicly available.
@@ II-A Differentiable manifolds
When the manifold $\mathcal{M}$ is also a Lie group, the tangent space at the identity, $T_{\mathcal{E}}\mathcal{M}$, is provided with a multiplication operation—typically non-associative—called the Lie bracket, which gives $T_{\mathcal{E}}\mathcal{M}$ an additional algebraic structure known as the Lie algebra, denoted $\mathfrak{m}$. This structure encapsulates the structure of the manifold. The elements of the tangent space can be interpreted as velocities or time derivatives, as well as error and increment variables used in estimation and control, for elements evolving on $\mathcal{M}$ [17].
@@
We now introduce the operators $\oplus$ and $\ominus$, which enable the combination of linear increments, expressed in the tangent vector space, with elements of the manifold. Following the convention in [17], increments expressed in the local frame (e.g., IMU readings) are always placed on the right-hand side of the composition,
@@
For a more exhaustive introduction to Lie theory, we refer the reader to [17].
II-B IKFoM
In the IESKF—and ESKF—formulation, the true-state is expressed as a suitable composition (e.g., linear sum, matrix product) of the nominal- and the error-state. The idea is to consider the nominal-state integrable in a non-linear fashion, and the error-state as linearly integrable and suitable for linear-Gaussian filtering [18]. Ultimately, the goal is to determine the linearized dynamics of the error-state.
@@ II-B IKFoM
When the manifold $\mathcal{M}$ is not a Lie group, the increment does not necessarily lie in its tangent space. Hence, we adopt the operator [^1] The $\oplus\ominus$ and $\boxplus$ notations are flipped w.r.t. IKFoM to be consistent with the notation in Solà et al. [17].
@@
In this work, the gravity vector is represented on the 2-sphere manifold $S^{2}(r)\triangleq\{\,\mathbf{x}\in\mathbb{R}^{3}:\|\mathbf{x}\|=r,\,r>0\,\}$, following the parameterization in [11]. The $S^{2}$ manifold defines a smooth surface embedded in $\mathbb{R}^{3}$ with constant norm $r$, enforcing a constant-norm directional constraint.
@@
The definitions of the $\oplus$ and $\ominus$ operators for $S^{2}$ are provided in Appendix V.
@@
The nominal-state $\mathbf{x}$ does not take into account the noise terms $\mathbf{w}$ in its dynamics, and other possible model nuances. As a consequence $\mathbf{x}$ accumulates error, which is collected in the error-state $\delta\mathbf{x}$, and estimated through the IESKF [18]. The true-state is defined as
@@
$\mathbf{F}_{\delta\mathbf{x}}$ $=\left.\frac{\partial\delta\mathbf{x}_{k+1}}{\partial\delta\mathbf{x}}\right|_{\delta\mathbf{x}=\mathbf{0},\mathbf{w}=\mathbf{0}},$ (7a)
@@
where Ad is the Adjoint map [17, eq. (30)], and $\mathbf{J}_{r}$ the right Jacobian [17, eq. (67)].
@@ II-C SGal(3)
The Special Galilean group $\mathrm{SGal}(3)$ provides a natural framework for the propagation of motion estimates over time, particularly in situations where temporal uncertainty plays a significant role [15]. By representing the state as a Galilean frame, time is intrinsically incorporated into the group structure, allowing spatial and temporal quantities to evolve jointly. This coupling captures the fundamental relationship between uncertainty in space and uncertainty in time, since uncertainty about when directly translates into uncertainty about where [13].
First used in robotics in [8] under the name of IMU deltas matrix Lie group, $\mathrm{SGal}(3)$ is a 10-dimensional Lie group that can be interpreted as the group of transformations that move points in both space and time—unlike $\mathrm{SE}(3)$ that only represents rigid motions in 3D space.
@@
For the definition and derivation of the $\exp$ and $\log$ maps associated with $\mathrm{SGal}(3)$, we refer the reader to [14].
@@ III-A System Overview
The overall pipeline is illustrated in Fig. 2. Incoming LiDAR scans are first deskewed using the IMU-integrated states, following the per-point deskewing strategy of [20], but with the motion model formulated on $\mathrm{SGal}(3)$ rather than on $\mathrm{SO}(3)\times\mathbb{R}^{6}$. Between LiDAR messages, raw IMU data is continuously propagated and a window of intermediate poses is stored so that each LiDAR point can be associated with its corresponding pose during deskewing.
@@ III-C Prediction
where $\hat{\mathbf{x}}$ is the propagated state, $\hat{\mathbf{P}}$ the propagated covariance, and $\mathbf{Q}$ is the covariance matrix of the noise $\mathbf{w}$ (see [18, eq. (262)–(265), (271)]). Between LiDAR scans, multiple IMU measurements are integrated sequentially. At the first IMU of a given interval, the propagation increment $\delta\mathbf{x}$ is initialized with the last LiDAR-updated state estimate, $\bar{\mathbf{x}}_{k-1}$.
@@ III-D Iterated Update
The iterations within the update step are initialized using the predicted prior, i.e., $\hat{\mathbf{x}}_{j}=\hat{\mathbf{x}}$ for $j=0$. At each iteration, for every downsampled point $\mathbf{p}_{i}$, we perform a kk–NN search in the i-Octree map in order to fit a local plane through them, yielding the point-to-plane constraint used in [20].
@@
The propagated state $\hat{\mathbf{x}}_{j}$ and covariance $\hat{\mathbf{P}}$ from (15b) impose a prior Gaussian distribution for the unkown true state $\mathbf{x}^{\star}_{j}$. More precisely, $\hat{\mathbf{P}}$ respresents the covariance of the following error state
@@
For a regular Lie group, $\mathbf{J}_{j}$ coincides with the inverse of the right Jacobian [17, eq. (83)], and it transports the propagated covariance matrix $\hat{\mathbf{P}}$ to the tangent space at $\hat{\mathbf{x}}_{j}$, at each iteration.
Following the iterated update formulation in [20], we gather all residuals into $\mathbf{z}_{j}$ and stack the partial derivatives of (17) into $\mathbf{H}_{j}$, then define the projected prior covariance as $\hat{\mathbf{P}}_{j}=(\mathbf{J}_{j})^{-1}\hat{\mathbf{P}}(\mathbf{J}_{j})^{-\top}$, and solve the Maximum A Posteriori (MAP) step via
@@
As in the i-Octree proposed in [24], each node can have up to eight children. This yields a shallower tree compared to a binary structure, and even in the worst-case branching factor, the effective complexity per level remains significantly lower. Moreover, children are indexed using Morton codes during subdivision, speeding up traversal and query operations.
@@ IV Evaluation
We evaluate LIMOncello (LIMO) on several public real-world datasets, including MCD [16], GrandTour [9], R-Campus [4], and a specific use case from the City Datasets [12]. LIMO is compared against RESPLE in its LIO configuration (R-LIO) and FAST-LIO2 (F-LIO2). All experiments are executed on a laptop equipped with an Intel Ultra 5 125H CPU and 32 GB RAM.
We interpolate trajectory estimates at ground truth timestamps and compute the RMSE of the absolute position error (APE) using the evo library [10]. Failures are denoted by ✗, and the best and second-best results are highlighted in bold and underline, respectively.
@@ IV-B Tree Efficiency Analysis
As shown in Fig. 4, ikd–Tree exhibits substantially higher computational and memory overhead (avg. 248 ms, 10.7 GB), mainly due to its continuous rebalancing and point redistribution. In contrast, both octree variants maintain real-time performance, with average runtimes of 21.7 ms (i-Octree) and 14.6 ms (LIMO i-Octree), and stable memory footprints of approximately 100–120 MB. Our i-Octree implementaion achieves the lowest runtime while retaining compact memory growth.
@@ V Conclusion
Following [11], the $\oplus$ and $\ominus$ operators are defined in terms of a local tangent basis $\mathbf{B}(\mathbf{x})\in\mathbb{R}^{3\times 2}$, whose columns span the tangent plane at $\mathbf{x}$. Specifically, for a point $\mathbf{x}\in S^{2}$ and a tangent increment $\boldsymbol{\tau}\in\mathbb{R}^{2}$, the $\oplus$ and $\ominus$ operations are given by
@@
To construct $\mathbf{B}(\mathbf{x})$, we follow [11] and define a rotation that maps the canonical basis onto a frame whose third axis aligns with $\mathbf{x}$. Using the reference axis $\mathbf{e}_{3}$, let
@@
With this choice, increments $\boldsymbol{\tau}$ live in $\mathbb{R}^{2}$ as illustrated in Fig. 5. Closed-form Jacobians are given in [11, eq. (69), (70)]. However, when $\|\mathbf{x}\times\mathbf{y}\|\to 0$, a first-order approximation is required to compute $\partial(\mathbf{y}\ominus\mathbf{x})/\partial\mathbf{y}$.
First-order approximation of atan2
Assume $\mathbf{x}$ and $\mathbf{y}$ are unit vectors. Define $x=\mathbf{x}^{\top}\mathbf{y},\text{and}\;y=\|\mathbf{x}\times\mathbf{y}\|.$ When $\mathbf{y}$ approaches $\mathbf{x}$, we have $(x,y)\to(1,0)$ and $\theta=\text{atan2}(y,x)$. The first-order Taylor expansion of $\text{atan2}(y,x)$ around $(x,y)=(1,0)$ yields
$\text{atan2}(y,x)\simeq y\qquad(x\approx 1,\,y\approx 0).$ (23)
Applying this approximation to the $\ominus$ operator we get $\mathbf{y}\ominus\mathbf{x}\simeq\mathbf{B}(\mathbf{x})^{\top}(\mathbf{x}\times\mathbf{y}).$ Thus,
$\frac{\partial(\mathbf{y}\ominus\mathbf{x})}{\partial\mathbf{y}}\simeq\mathbf{B}(\mathbf{x})^{\top}\,[\mathbf{x}]_{\times}$ (24)
Antipodal singularity
When $\mathbf{x}$ and $\mathbf{y}$ are antiparallel, $\theta=\pi$, but the rotation axis is undefined, there exist infinitely many tangent vectors $\boldsymbol{\tau}$, with $\|\boldsymbol{\tau}\|=\pi$, that rotate $\mathbf{x}$ into $\mathbf{y}$. Following the convention in [11], we choose

### v2

# LIMOncello: Iterated Error-State Kalman Filter on the SGal(3) Manifold for Fast LiDAR–Inertial Odometry

@@ I Introduction
Light detection and ranging (LiDAR) sensors complement vision in such setups by providing long-range, illumination-invariant, and metrically accurate depth, which translates into reliable odometry and mapping across diverse environments. However, LiDAR introduces distinct challenges: i) processing a massive, continuous stream of data in a timely and efficient manner, ii) overcome the scarcity of reliable geometric features in structurally poor environments, and iii) correcting the scan deformation caused by sensor movement during data capture [FAST-LIO, FAST-LIO2].
@@
To address these issues, LiDAR–Inertial Odometry (LIO) combines LiDAR with high-rate inertial measurements. High-rate Inertial Measurement Unit (IMU) measurements enable per-point deskewing and cover the latency between LiDAR frames, while LiDAR provides metric structure to correct the robot’s pose. Most LIO systems are organized into a front-end and a back-end [Faster-LIO]. The front-end maintains a history map and performs fast data association. This process is based on classical point-cloud registration, which works by finding the nearest neighbors for each incoming point. These neighbors are then used to fit a simple local model (typically a plane or a line) and form the residuals for optimization.
@@
The back-end is responsible for state optimization, with the two common families being graph-based smoothing and filtering. Under limited onboard processing, filtering based on the Extended Kalman Filter and its iterated error-state (IESKF) variant is often preferred [IKFoM], as exemplified by FAST-LIO1 [FAST-LIO] and FAST-LIO2 works [FAST-LIO2].
@@
A key step forward was the IKFoM [IKFoM] framework, which formalized the IESKF on differentiable manifolds and packaged it into a practical toolkit. This gave a canonical discrete-time representation and clean separation between manifold operations and system-specific models.
@@
With a robust and reusable back-end in place, the community has focused effort on the front-end problems of features and computation: designing faster spatial data structures and residual evaluations that sustain real-time rates without sacrificing accuracy. The community has improved classical k-d tree map structures with incremental variants such as the ikd-Tree [ikdtree] and sparse voxel representations like iVox [Faster-LIO], enabling efficient real-time updates and queries. Building on these advances, recent methods have shifted from point-to-map registration to point-to-feature formulations, where points are aligned with local geometric primitives (e.g., planes or edges) rather than individual map points [VoxelMap++, C3P-VoxelMap].
@@
Alongside discrete-time pipelines, continuous-time motion models have gained attraction. By representing the trajectory with piecewise functions and evaluating residuals at each point timestamp, these methods improve deskewing and alignment during fast motion or asynchronous sensing. Recent systems such as DLIO [DLIO] adopt point-wise continuous-time motion compensation, and others like RESPLE [RESPLE] show that investing in state representation and motion modeling can yield competitive accuracy while preserving real-time performance.
@@
Most LIO frameworks rely on what is often referred to as a compound [IKFoM] or bundle manifold [microLie] representation of the state. In these formulations, the system’s pose, and velocity are defined on separate sub-manifolds—usually $\mathrm{SO}(3)\times\mathbb{R}^{6}$. While this decomposition simplifies the implementation, it also neglects certain kinematic couplings between translational and rotational motion, leading to an approximation that may lose accuracy under highly dynamic motion or under sustained low-observability conditions, e.g., near-zero acceleration that reduces IMU observability, or degenerate LiDAR environments such as long corridors.
@@
Beyond Euclidean and compound representations, several Lie-group manifolds have been adopted in robotics. $\mathrm{SE}(3)$ is widely used for pose estimation, while $\mathrm{SE}_{2}(3)$ became standard in Invariant EKF-based methods [iEKF], with extensions such as $\mathrm{SE}_{N}(3)$ [SI-LIO] and $\mathrm{SGal}(3)$ explored more recently in equivariant filtering [SGal_pre]. However, their use in tightly coupled LIO-SLAM systems remains limited.
@@
  • Efficient spatial data structure: We replace the widely used ikd-Tree with a refactored, lightweight i-Octree implementation [ioctree], achieving faster incremental updates and lower computational overhead by avoiding costly tree rebalancing.
  • Modular and extensible implementation: LIMOncello is implemented using the manif library [manif], resulting in a clean, modular codebase that is easily extensible and publicly available.
@@ II-A Differentiable manifolds
When the manifold $\mathcal{M}$ is also a Lie group, the tangent space at the identity, $T_{\mathcal{E}}\mathcal{M}$, is provided with a multiplication operation—typically non-associative—called the Lie bracket, which gives $T_{\mathcal{E}}\mathcal{M}$ an additional algebraic structure known as the Lie algebra, denoted $\mathfrak{m}$. This structure encapsulates the structure of the manifold. The elements of the tangent space can be interpreted as velocities or time derivatives, as well as error and increment variables used in estimation and control, for elements evolving on $\mathcal{M}$ [microLie].
@@
We now introduce the operators $\oplus$ and $\ominus$, which enable the combination of linear increments, expressed in the tangent vector space, with elements of the manifold. Following the convention in [microLie], increments expressed in the local frame (e.g., IMU readings) are always placed on the right-hand side of the composition,
@@
For a more exhaustive introduction to Lie theory, we refer the reader to [microLie].
II-B IESKF on Manifolds
In the IESKF—and ESKF—formulation, the true-state is expressed as a suitable composition (e.g., linear sum, matrix product) of the nominal- and the error-state. The idea is to consider the nominal-state integrable in a non-linear fashion, and the error-state as linearly integrable and suitable for linear-Gaussian filtering [quaternionESKF]. Ultimately, the goal is to determine the linearized dynamics of the error-state.
@@ II-B IESKF on Manifolds
When the manifold $\mathcal{M}$ is not a Lie group, the increment does not necessarily lie in its tangent space. Hence, we adopt the operator [^1] The $\oplus\ominus$ and $\boxplus$ notations are flipped w.r.t. IKFoM to be consistent with the notation in Solà et al. [microLie].
@@
In this work, the gravity vector is represented on the 2-sphere manifold $S^{2}(r)\triangleq\{\,\mathbf{x}\in\mathbb{R}^{3}:\|\mathbf{x}\|=r,\,r>0\,\}$, following the parameterization in [IKFoM]. The $S^{2}$ manifold defines a smooth surface embedded in $\mathbb{R}^{3}$ with constant norm $r$, enforcing a constant-norm directional constraint.
@@
The definitions of the $\oplus$ and $\ominus$ operators for $S^{2}$ are provided in Appendix VI.
@@
The nominal-state $\mathbf{x}$ does not take into account the noise terms $\mathbf{w}$ in its dynamics, and other possible model nuances. As a consequence $\mathbf{x}$ accumulates error, which is collected in the error-state $\delta\mathbf{x}$, and estimated through the IESKF [quaternionESKF]. The true-state is defined as
@@
$\mathbf{F}_{\delta\mathbf{x}}$ $=\left.\frac{\partial\delta\mathbf{x}_{k+1}}{\partial\delta\mathbf{x}_{k}}\right|_{\delta\mathbf{x}=\mathbf{0},\mathbf{w}=\mathbf{0}},$ (7a)
@@
where Ad is the Adjoint map [microLie, eq. (30)], and $\mathbf{J}_{r}$ the right Jacobian [microLie, eq. (67)].
@@ II-C SGal(3)
The Special Galilean group $\mathrm{SGal}(3)$ provides a natural framework for the propagation of motion estimates over time, particularly in situations where temporal uncertainty plays a significant role [mahony2025galileansymmetryrobotics]. By representing the state as a Galilean frame, time is intrinsically incorporated into the group structure, allowing spatial and temporal quantities to evolve jointly. This coupling captures the fundamental relationship between uncertainty in space and uncertainty in time, since uncertainty about when directly translates into uncertainty about where [makingspacetimespecial].
First used in robotics in [IMUDeltas] under the name of IMU deltas matrix Lie group, $\mathrm{SGal}(3)$ is a 10-dimensional Lie group that can be interpreted as the group of transformations that move points in both space and time—unlike $\mathrm{SE}(3)$ that only represents rigid motions in 3D space.
@@
For the definition and derivation of the $\exp$ and $\log$ maps associated with $\mathrm{SGal}(3)$, we refer the reader to [kelly2025galileangroupsgal3].
@@ III-A System Overview
The overall pipeline is illustrated in Fig. 2. Incoming LiDAR scans are first deskewed using the IMU-integrated states, following the per-point deskewing strategy of [FAST-LIO2], but with the motion model formulated on $\mathrm{SGal}(3)$ rather than on $\mathrm{SO}(3)\times\mathbb{R}^{6}$. Between LiDAR messages, raw IMU data is continuously propagated and a window of intermediate poses is stored so that each LiDAR point can be associated with its corresponding pose during deskewing.
@@ III-C Prediction
where $\hat{\mathbf{x}}$ is the propagated state, $\hat{\mathbf{P}}$ the propagated covariance, and $\mathbf{Q}$ is the covariance matrix of the noise $\mathbf{w}$ (see [quaternionESKF, eq. (262)–(265), (271)]). Between LiDAR scans, multiple IMU measurements are integrated sequentially. At the first IMU of a given interval, $\hat{\mathbf{x}}$ is initialized with the last LiDAR-updated state estimate, $\bar{\mathbf{x}}_{k-1}$.
@@ III-D Iterated Update
The iterations within the update step are initialized using the predicted prior, i.e., $\hat{\mathbf{x}}_{j}=\hat{\mathbf{x}}$ for $j=0$. At each iteration, for every downsampled point $\mathbf{p}_{i}$, we perform a kk–NN search in the i-Octree map in order to fit a local plane through them, yielding the point-to-plane constraint used in [FAST-LIO2].
@@
The propagated state $\hat{\mathbf{x}}_{j}$ and covariance $\hat{\mathbf{P}}$ from (15b) impose a prior Gaussian distribution for the unknown true state $\mathbf{x}^{\star}_{j}$. More precisely, $\hat{\mathbf{P}}$ represents the covariance of the following error state
@@
For a regular Lie group, $\mathbf{J}_{j}$ coincides with the inverse of the right Jacobian [microLie, eq. (83)], and it transports the propagated covariance matrix $\hat{\mathbf{P}}$ to the tangent space at $\hat{\mathbf{x}}_{j}$, at each iteration.
Following the iterated update formulation in [FAST-LIO2], we gather all residuals into $\mathbf{z}_{j}$ and stack the partial derivatives of (17) into $\mathbf{H}_{j}$, then define the projected prior covariance as $\hat{\mathbf{P}}_{j}=(\mathbf{J}_{j})^{-1}\hat{\mathbf{P}}(\mathbf{J}_{j})^{-\top}$, and solve the Maximum A Posteriori (MAP) step via
@@
As in the i-Octree proposed in [ioctree], each node can have up to eight children. This yields a shallower tree compared to a binary structure, and even in the worst-case branching factor, the effective complexity per level remains significantly lower. Moreover, children are indexed using Morton codes during subdivision, speeding up traversal and query operations.
@@ IV Evaluation
We evaluate LIMOncello (LIMO) on several public real-world datasets, including MCD [MCDDataset], GrandTour [GrandTour], R-Campus [RESPLE], and a specific use case from the City Datasets [MA-LIO]. LIMO is compared against RESPLE in its LIO configuration (R-LIO) and FAST-LIO2 (F-LIO2). All experiments are executed on a laptop equipped with an Intel Ultra 5 125H CPU and 32 GB RAM.
We interpolate trajectory estimates at ground truth timestamps and compute the RMSE of the absolute position error (APE) using the evo library [grupp2017evo]. Failures are denoted by ✗, and the best and second-best results are highlighted in bold and underline, respectively.
@@ IV-B Tree Efficiency Analysis
As shown in Fig. 4, ikd–Tree exhibits substantially higher computational and memory overhead (avg. 248 ms, 10.7 GB), mainly due to its continuous rebalancing and point redistribution. In contrast, both octree variants maintain real-time performance, with average runtimes of 21.7 ms (i-Octree) and 14.6 ms (LIMO i-Octree), and stable memory footprints of approximately 100–120 MB. Our i-Octree implementation achieves the lowest runtime while retaining compact memory growth.
@@ VI Conclusion
Following [IKFoM], the $\oplus$ and $\ominus$ operators are defined in terms of a local tangent basis $\mathbf{B}(\mathbf{x})\in\mathbb{R}^{3\times 2}$, whose columns span the tangent plane at $\mathbf{x}$. Specifically, for a point $\mathbf{x}\in S^{2}$ and a tangent increment $\boldsymbol{\tau}\in\mathbb{R}^{2}$, the $\oplus$ and $\ominus$ operations are given by
@@
To construct $\mathbf{B}(\mathbf{x})$, we follow [IKFoM] and define a rotation that maps the canonical basis onto a frame whose third axis aligns with $\mathbf{x}$. Using the reference axis $\mathbf{e}_{3}$, let
@@
With this choice, increments $\boldsymbol{\tau}$ live in $\mathbb{R}^{2}$ as illustrated in Fig. 5. Closed-form Jacobians are given in [IKFoM, eq. (69), (70)].
However, when $\theta\to 0$, a first-order Taylor approximation is applied to atan2, and (21b) simplifies to $\mathbf{y}\ominus\mathbf{x}\simeq\mathbf{B}(\mathbf{x})^{\top}(\mathbf{x}\times\mathbf{y}).$ Therefore,
$\frac{\partial(\mathbf{y}\ominus\mathbf{x})}{\partial\,\mathbf{y}}\simeq\mathbf{B}(\mathbf{x})^{\top}\,[\mathbf{x}]_{\times}.$ (23)
Conversely, when $\theta\to\pi$, there exist infinitely many tangent vectors that rotate $\mathbf{x}$ into $\mathbf{y}$. Following the convention in [IKFoM], we choose